OpenStudy (anonymous):

Sales tax in Pennsylvania is 6%. Create an equation for the total price (cost plus tax) of a purchase in PA in terms of its cost. How much would a person pay for a car whose cost is $32,000? What's the equation?

OpenStudy (jackellyn):

6/100=0.06 32,000*0.06=tax 32,000+tax=cost or Total price = item price*(1.06) Total price = 32000*(1.06) = 33920